Find a number such that 8 times the number of 36 greater than the opposite of the number.
Answer by ewatrrr(24785)


Hi,
Let x be the number
writing as question reads
8x = -x + 36
9x = 36
x = 4
checking our answer
8*4 = -4 + 36 = 32